OpenStudy (anonymous):

What is the answer of an equation of a straight line through the point (-1,-7) and parallel to y=2x+3. (Show in working steps please.)

OpenStudy (anonymous):

The parallel line will have the same slope as the other line, m=2. You can use point-slope formula for this.

OpenStudy (anonymous):

Could you please show me this answer in working steps I still dont understand..

OpenStudy (anonymous):

Do you know point-slope form of an equation?

OpenStudy (anonymous):

Not really, no.

OpenStudy (anonymous):

It looks like this \[\large y-y_1=m(x-x_1)\] It is derived from the slope formula \[\large m=\frac{y-y_1}{x-x_1}\]

OpenStudy (anonymous):

Just plug in the given info, m=2, x1=-1, y1=-7.

OpenStudy (anonymous):

Then, if you want, you can rearrange it into slope-intercept form (y=mx+b) if you prefer.

OpenStudy (anonymous):

Would you please do it for me I still can't get the right answer..

OpenStudy (anonymous):

Show me what you did get, and I'll point out any mistakes.

OpenStudy (anonymous):

Nothing... I dont kno how to start it off!

OpenStudy (anonymous):

Could you help me get started please? In my question format..